# An $\ell^{1}$-Plug-and-Play Approach for Magnetic Particle Imaging Using a Zero Shot Denoiser with Validation on the 3D Open MPI Dataset

### v1

@@ Abstract
Magnetic particle imaging (MPI) is an emerging medical imaging modality which has gained increasing interest in recent years. Among the benefits of MPI are its high temporal resolution, and that the technique does not expose the specimen to any kind of ionizing radiation. It is based on the non-linear response of magnetic nanoparticles to an applied magnetic field. From the electric signal measured in receive coils, the particle concentration has to be reconstructed. Due to the ill-posedness of the reconstruction problem, various regularization methods have been proposed for reconstruction ranging from early stopping methods, via classical Tikhonov regularization and iterative methods to modern machine learning approaches. In this work, we contribute to the latter class: we propose a plug-and-play approach based on a generic zero-shot denoiser with an $\ell^{1}$-prior. Moreover, we develop parameter selection strategies. Finally, we quantitatively and qualitatively evaluate the proposed algorithmic scheme on the 3D Open MPI data set with different levels of preprocessing.
@@ 1 Introduction
In brief, the principle and the procedure behind MPI are as follows: (i) the specimen to be scanned is injected with a tracer containing the magnetic nanoparticles (MNs); (ii) a dynamic magnetic field with a low-field volume (LFV) is applied and the LFV is spatially moved; (iii) the LFV acts as a sensitive region whose motion causes the particles to induce a voltage in the receive coils. The imaging task of MPI is the reconstruction of the spatial distribution of the particles in the tracer from the voltage signal acquired during a scan. There are currently two main types of approaches to reconstructing particle distributions: measurement-based approaches and model-based approaches. Measurement-based approaches typically acquire a system matrix: for each column, one performs scanning of a delta probe located in a corresponding voxel of a chosen 3D voxel grid [35, 65, 48, 41]. More precisely, a 3D grid of the volume to be scanned inside the scanner’s bore is considered; a probe with a reference concentration of tracer is iteratively positioned and scanned at each voxel of the grid. This way the response of the scanning system to discrete delta impulses at each voxel position is collected and stored in the system matrix $A.$ Assuming linearity, the acquired signal $f$ of the scan of a specimen is then obtained as a superpositon of the scans of the delta impulses. For reconstruction, one has to solve a corresponding system of equations $Au=f,$ where the symbol $u$ denotes the desired concentration; loosely speaking, one has to invert the system matrix for the measured signal $f$. Due to the presence of noise and the ill-conditioning of the system matrix (e.g., [38, 56]), regularization techniques are needed for the inversion as discussed later. The development of model-based reconstruction techniques is another direction of research aiming at avoiding or significantly reducing the calibration procedures needed for measuring the system matrix; for details, cf. for instance  [47, 51, 18, 19, 21, 45, 9]. In this work we consider the measurement approach and focus on regularized machine-learning based reconstruction in that scenario.
@@
Due to the ill-conditioning of the system matrix and the related ill-posedness of the reconstruction problem in MPI [34, 45, 33] small perturbations in the data can lead to significant errors of the reconstruction; thus, regularization is needed [7]. Classical regularization approaches in MPI include Tikhonov regularization [66, 38] using conjugate gradients and the Kaczmarz method. To improve noise suppression capabilities and further features, structural priors based on the $\ell^{1}$-norm have been used within iterative schemes; e.g.[56]. For a broader overview we refer to [13].
In recent years, there is increasing interest in reconstruction methods involving machine learning approaches. Among these are methods learning the reconstruction by directly training a neural network, e.g. [27, 63]. Further, reconstruction based on the Deep Image Prior (DIP) has been proposed [14, 61]. Recently, Deep Equilibrium Models have been employed in the context of MPI [24, 5]. Deep learning approaches have been further used to improve the spatial resolution of the reconstructions [53], and to reduce the calibration time, for example by partially acquiring system matrix on a down-sampled grid and use deep-learning for super-resolution of the system matrix [22, 52, 23, 68].
A particular class of machine learning approaches are plug and play (PnP) approaches. They are inspired by iterative schemes realizing energy minimization: building on an iterative scheme, they replace the denoiser implied by the scheme by an arbitrary one; instead of classical denoising, typically machine-learning based denoisers are employed [70]. PnP approaches have been used for a variety of imaging modalitites, e.g., [2, 55] as well as for computer vision tasks, e.g., [70, 57]. These approaches are conservative in the sense that machine learning only enters the scheme via iterative denoising and that well-known and well-understood components from iterative schemes and classical Tikhonov regularization are still integrated into the reconstruction scheme.
@@
In this work we develop a plug and play (PnP) approach for MPI reconstruction based on a generic (zero-shot) denoiser, a half quadratic splitting scheme plus an $\ell^{1}$-prior. Moreover, we develop parameter choice strategies, and evaluate the proposed method on the publicly available Open MPI data set. In detail, we make the following contributions:
@@
    In the proposed $\ell^{1}$-PnP method we employ a generic (zero-shot) image denoiser. This means the denoiser is not transfer learned on MPI or similar MRI data. (To apply the 2D image denoiser on a 3D volume, we use slicing.) The approach is inspired by the recent observations in Computer Vision that larger models pretrained on larger training sets perform extremely well even when employed in off-distribution contexts. We further employ a shrinkage/ $\ell^{1}$-prior which is motivated by the works [56, 32] which use a corresponding prior in an energy minimization frameworks in MPI reconstruction.
@@
    We propose parameter selection strategies.
@@
    We quantitatively and qualitatively evaluate our approach on the full 3D Open MPI data set [39] which presently represents the only publicly available real MPI data set. In particular, we quantitatively and qualitatively compare our method with the DIP approach [14] using the preprocessing pipeline [31] and classical Tikhonov regularization. Further, we evaluate the approach with the parameter selection and stopping criteria strategies. Moreover, we investigate the influence of the preprocessing on the reconstruction quality.
Quantitative and qualitative evaluation of a PnP approach (including comparison with another machine learning approach) on the full 3D Open MPI data set has not been done before. (We note that quantitative evaluation is possible on the Open MPI dataset since a reference CAD model is available in the dataset and serves the purpose of a ground truth [39]). Further, employing a zero-shot denoiser [70] (trained on a large set of natural images) is new.
@@
In Section 2 we develop the method: we describe the formal MPI reconstruction problem (Section 2.1), the preprocessing pipeline of the MPI data (Section 2.2), we present the proposed $\ell^{1}$-PnP approach (Section 2.3), and we consider parameter selection (including a stopping criterion) in Section 2.4. In Section 3 we quantitatively and qualitatively evaluate the proposed $\ell^{1}$-PnP algorithm on the OpenMPI Dataset: we compare with the DIP and Tikhonov regularization (Section 3.3), investigate the influence of different levels of preprocessing (Section 3.4) and consider a setup without preprocessing in Section 3.5. We conclude with Section 4 which contains a discussion and a conclusion part.
@@ 2 Method
We first recall the formal problem of MPI reconstruction based on a system matrix approach in Section 2.1. Then we discuss the employed preprocessing in Section 2.2. Finally, we derive the proposed $\ell^{1}$-PnP approach using a zero-shot denoiser in Section 2.3 and consider parameter selection in Section 2.4.
2.1 Reconstruction Problem in MPI Using a System Matrix
@@
The task of MPI reconstruction is to determine the density $u$ of iron oxide nanoparticles contained in an object. We briefly describe the basic principle of a field-free point (FFP) scanner; for details we refer to [37]. A specific static magnetic field is applied to saturate the nanoparticles everywhere in a field of view but in a small region. Such low-field region is a small neighborhood of a point called the Field-Free Point (FFP). A second dynamic magnetic field is applied to steer the FFP in the field of view. The movement of the FFP constitutes a scan: the change of magnetization of the iron oxide particles due to the movement of the FFP induces a voltage in a set of receive coils which constitutes the measured data $f$. The relation between the concentration $u$ and the measured data $f$ is modeled as a linear mapping called the system function or system matrix $A$. In order to determine the system matrix/function both model and measurement-based approaches have been applied; cf. [48, 36] for instance. In real data scenarios, at present, the measurement-based approach is mostly used: the system matrix is obtained by measuring the system response to delta probes centered at all voxel. We consider MPI reconstruction in three dimesional space using real measurement data from the Open MPI dataset [39]. Hence, the reconstructed 3D image/volume $\tilde{u}$ lives in the space $X=\mathbb{R}^{n_{1}\times n_{2}\times n_{3}},$ where $n_{1}=n_{2}=n_{3}=19$ for the Open MPI data set. We denote the total number of voxels by $N=n_{1}n_{2}n_{3}=19^{3},$ and the number of measurements by $M$; here $M$ depends on the preprocessing applied to the MPI data (cf. Section 2.2). In particular, a scan in the Open MPI dataset contains the data coming from three different channels and the calibration data is Fourier transformed. This means that for each of the three channels the real and imaginary parts of the Fourier coefficients are considered and stacked, yielding a real-valued submatrix $A_{l}$. Finally, the full matrix $A$ is formed by stacking the submatrices $A_{l}$ for each channel $l$ (for a detailed explanation we refer the interested reader to [31]). In order to avoid notation issues, we use the following conventions: we use the symbol $u\in\mathbb{R}^{N}$ to denote the (columnwise) vectorization of the volume $\tilde{u}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}.$ We then identify the system matrix with a matrix $\tilde{A}\in\mathbb{R}^{M\times N}$ such that $\tilde{A}\tilde{u}=Au$ for all volumes $\tilde{u}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$. We further identify the adjoint $A^{\ast}$ of $A$ with the matrix $\tilde{A}^{T}$; in particular, for any $g\in\mathbb{R}^{M},$ we have that $A^{\ast}g$ equals the reshape operation (which inverts the vectorization operation) of $\tilde{A}^{T}g$. Using this notation, we have the relation
@@
where $\tilde{A}\in\mathbb{R}^{M\times N}$ denotes the system matrix, $u$ denotes the discrete nano-particle density/volume, $f$ the measured data, and the symbol $\eta$ denotes the noise on the measurements. In particular, the $N=19^{3}$ columns of $\tilde{A}$ are obtained by measuring the scanner’s response to delta probes sitting in the respective voxels. The reconstruction task consists of determining $u$ in Eq. (1) given measurements $f$ (and $A$). This is challenging due to noise and the ill-conditioning of $A$ and requires regularization as already discussed.
2.2 Preprocessing of MPI Data
We briefly discuss the preprocessing of the Open MPI data employed. It is inspired but not identical to the preprocessing proposed of [31]. For this reason, and since we discuss our scheme w.r.t. different levels of preprocessing in the experimental section, we briefly explain the used method.
Preprocessing is applied to both calibration and phantom scan data. Since the calibration data are stored in the Fourier domain we transform the scan data (given as a time series) to the Fourier domain. We further split the real and imaginary part of the Fourier coefficients for both calibration and phantom data. Although any preprocessing step may decrease the number of rows of $A,$ we always use the same symbol $M$ to denote these numbers.
@@ 2.2 Preprocessing of MPI Data
Due to analog filtering, the measurement data show a very high variance in the frequencies up to 75kHz. We therefore exclude frequencies lower than 80kHz. Further frequency selection avoiding frequencies with systematic strong noise or with low signal-to-noise ratio (SNR) has been considered in [31]. To reduce the number of variables at play and to avoid estimations of the SNR we consider no SNR-based thresholding in the preprocessing here.
@@
The noisy measured phantom data $f$ is of the form $f=f_{\mathrm{cl.}}+b_{f}+\eta_{f}$, with clean signal $f_{\mathrm{cl.}}$ containing the information on $u,$ the background signal $b_{f}$ produced by the drive field and perturbations in the measurement chain, and noise $\eta_{f}.$ Accordingly, we decompose the system matrix as $A=A_{\mathrm{cl}}+b_{A}+\eta_{A},$ with clean signals $A_{\mathrm{cl}},$ background $b_{A}$ and noise $\eta_{A}.$ To exclude the background signals $b_{f},$ $b_{A}$, background measurements with no phantom in the scanner are performed after the scan and during the calibration procedure. To account for the background signal to be dynamic, it is repeated after every $19$th delta probe scan (after running through the $x$-axis) and a corresponding convex combination is subtracted. Each phantom and delta scan is performed $1000$ times and the average is taken as input for the reconstruction. (By linearity, the averaging and subtraction is in fact carried out in the Fourier domain.) For a more precise description we refer to the Sec. S1 in the supplementary material.
@@
The low rank approximation is obtained by selecting $K\leq\min\{M,N\}$ and by performing a randomized Singular Value Decomposition (rSVD) [31, 26], yielding $(\tilde{U}_{K},\tilde{\Sigma}_{K},\tilde{V}_{K})$ and considering for the reconstruction the transformed matrix $\tilde{U}_{K}^{\ast}A\in\mathbb{R}^{K\times N}$ and the transformed data $\tilde{f}=\tilde{U}_{K}^{\ast}f\in\mathbb{R}^{K}$. Because we aim at using the rSVD obtained we store $\tilde{V}_{K}$ and $\mathrm{diag}(\tilde{\Sigma}_{K})\in\mathbb{R}^{K}$ in place of $\tilde{U}_{K}^{\ast}A$.
2.3 The Proposed Plug-and-Play Approach with a Zero-Shot Denoiser
We now describe the proposed $\ell^{1}$-PnP approach. The regularized inversion of the ill-posed MPI reconstruction problem in Eq. (1) can be formulated as a minimization problem of the form

### v2

# An $\ell^{1}$-Plug-and-Play Approach for MPI Using a Zero Shot Denoiser with Evaluation on the 3D Open MPI Dataset

@@ Abstract
Objective: Magnetic particle imaging (MPI) is an emerging medical imaging modality which has gained increasing interest in recent years. Among the benefits of MPI are its high temporal resolution, and that the technique does not expose the specimen to any kind of ionizing radiation. It is based on the non-linear response of magnetic nanoparticles to an applied magnetic field. From the electric signal measured in receive coils, the particle concentration has to be reconstructed. Due to the ill-posedness of the reconstruction problem, various regularization methods have been proposed for reconstruction ranging from early stopping methods, via classical Tikhonov regularization and iterative methods to modern machine learning approaches. In this work, we contribute to the latter class: we propose a plug-and-play approach based on a generic zero-shot denoiser with an $\ell^{1}$-prior.
Approach: We validate the reconstruction parameters of the method on a hybrid dataset and compare it with the baseline Tikhonov, DIP and the previous PP-MPI, which is a plug-and-play method with denoiser trained on MPI-friendly data.
Main results: We offer a quantitative and qualitative evaluation of the zero-shot plug-and-play approach on the 3D Open MPI dataset. Moreover, we show the quality of the approach with different levels of preprocessing of the data.
Significance: The proposed method employs a zero-shot denoiser which has not been trained for the MPI task and therefore saves the cost for training. Moreover, it offers a method that can be potentially applied in future MPI contexts.
@@ 1 Introduction
In brief, the principle and the procedure behind MPI are as follows: (i) the specimen to be scanned is injected with a tracer containing the magnetic nanoparticles (MNs) and placed in the scanner; (ii) a dynamic magnetic field with a low-field volume (LFV) is applied and the LFV is spatially moved; (iii) the LFV acts as a sensitive region whose motion causes the particles to induce a voltage in the receive coils. The imaging task of MPI is the reconstruction of the spatial distribution of the particles in the tracer from the voltage signal acquired during a scan. There are currently two main types of approaches to reconstructing particle distributions: measurement-based approaches and model-based approaches. Measurement-based approaches typically acquire a system matrix: for each column of the matrix, one performs scanning of a delta probe located in a corresponding voxel of a chosen 3D voxel grid [35, 65, 48, 41]. More precisely, a 3D grid of the volume to be scanned inside the scanner’s bore is considered; a probe with a reference concentration of tracer is iteratively positioned and scanned at each voxel of the grid. This way the response of the scanning system to discrete delta impulses at each voxel position is collected and stored in the system matrix $A$. Assuming linearity, the acquired signal $f$ of the scan of a specimen is then obtained as a superpositon of the scans of the delta impulses. For reconstruction, one has to solve a corresponding system of equations $Au=f,$ where the symbol $u$ denotes the desired concentration; loosely speaking, one has to invert the system matrix for the measured signal $f$. Due to the presence of noise and the ill-conditioning of the system matrix (e.g., [38, 56]), regularization techniques are needed for the inversion as discussed later. The development of model-based reconstruction techniques is a direction of research aiming at avoiding or significantly reducing the calibration procedures needed for measuring the system matrix; for details, cf. for instance  [47, 51, 18, 19, 21, 45, 9]. In this work we consider the measurement approach and focus on regularized machine learning-based reconstruction in that scenario.
@@
Due to the ill-conditioning of the system matrix and the related ill-posedness of the reconstruction problem in MPI [34, 45, 33], small perturbations in the data can lead to significant errors in the reconstruction; thus, regularization is needed [7]. Classical regularization approaches in MPI include Tikhonov regularization [66, 38] using conjugate gradients and the Kaczmarz method. To improve noise suppression capabilities and further features, structural priors based on the $\ell^{1}$-norm have been used within iterative schemes; e.g.[56]. For a broader overview we refer to [13].
In recent years, there is increasing interest in reconstruction methods involving machine learning approaches. Among these are methods learning the reconstruction by training a neural network, e.g. [27, 63]. Further, reconstruction based on the Deep Image Prior (DIP) has been proposed [14, 61]. Recently, Deep Equilibrium Models have been employed in the context of MPI [24, 5]. Deep learning approaches have been further used to improve the spatial resolution of the reconstructions [53], and to reduce the calibration time, for example by partially acquiring a system matrix on a down-sampled grid and use deep-learning for super-resolution of the system matrix [22, 52, 23, 68].
A particular class of machine learning-based approaches are Plug and Play (PnP) approaches. They are inspired by iterative schemes realizing energy minimization: building on a splitting scheme, they replace the denoiser implied by the slitting with an arbitrary one; instead of classical denoising, typically machine learning-based denoisers are employed [70]. PnP approaches have been used for a variety of imaging modalitites, e.g., [2, 55] as well as for computer vision tasks, e.g., [70, 57]. These approaches are conservative in the sense that machine learning only enters the scheme via iterative denoising and that well-known and well-understood components from iterative schemes and classical Tikhonov regularization are still integrated into the reconstruction scheme.
@@
In this work we consider a plug and play (PnP) approach for MPI reconstruction based on a generic (zero-shot) denoiser, a half quadratic splitting scheme plus an additional $\ell^{1}$-prior. Moreover, we validate the parameters on a hybrid dataset and evaluate the proposed methods on the publicly available Open MPI dataset. In detail, we make the following contributions:
@@
    In the proposed PnP method we employ a generic (zero-shot) image denoiser. This means the denoiser is not transfer learned on MPI or similar MRI data. (To apply the 2D image denoiser on a 3D volume, we use slicing.) The approach is inspired by the recent observations in Computer Vision that larger models pretrained on larger training sets perform well even when employed in off-distribution contexts. We further employ a shrinkage/$\ell^{1}$-prior which is motivated by the works [56, 32] which use a corresponding prior in an energy minimization framework in MPI reconstruction.
@@
    We validate the parameters on a hybrid 3D MPI dataset.
@@
    We quantitatively and qualitatively evaluate our approach on the full 3D Open MPI dataset [39] which presently represents the only publicly available real MPI dataset. In particular, we quantitatively and qualitatively compare our method with the DIP approach [14] using the preprocessing pipeline in [31], with the classical Tikhonov regularization and previous PnP approaches [4]. Moreover, we investigate the influence of the preprocessing on the reconstruction quality.
Both quantitative and qualitative evaluation of a PnP approach (including comparison with another machine learning approach) on the full 3D Open MPI dataset has not been done before. (We note that quantitative evaluation is possible on the Open MPI dataset since a reference CAD model is available in the dataset and serves the purpose of a ground truth [39]). Further, employing a zero-shot denoiser [70] (trained on a large set of natural images) in the context of MPI is new.
@@
In Section 2 we develop the method: we describe the formal MPI reconstruction problem (Section 2.1), the preprocessing pipeline of the MPI data (Section 2.2), we present the proposed ZeroShot-PnP approach with and without the additional $\ell^{1}$-prior (Section 2.3), we consider parameter selection in Section 2.4 and describe the hybrid dataset created for the validation of the parameters in Section 2.5. In Section 3 we quantitatively and qualitatively evaluate the proposed ZeroShot-$\ell^{1}$-PnP algorithm on the OpenMPI Dataset: we compare with the DIP, Tikhonov regularization and previous Plug-and-Play approaches (Section 3.3 and 3.4), investigate the influence of different levels of preprocessing (Section 3.5) and consider a setup without preprocessing in Section 3.6. We conclude with Section 4 which contains a discussion and a conclusion part.
@@ 2 Method
We first recall the formal problem of MPI reconstruction based on a system matrix approach in Section 2.1. Then we discuss the employed preprocessing in Section 2.2. Finally, we derive the proposed ZeroShot-$\ell^{1}$-PnP approach using a zero-shot denoiser in Section 2.3 and consider parameter selection in Section 2.4.
2.1 The Reconstruction Problem in MPI Using a System Matrix
@@
The task of MPI reconstruction is to determine the concentration $u$ of iron oxide nanoparticles contained in an object. We briefly describe the basic principle of a field-free point (FFP) scanner; for details we refer to [37]. A specific static magnetic field is applied to saturate the nanoparticles everywhere in a field of view but in a small region. Such low-field region is a small neighborhood of a point called the Field-Free Point (FFP). A second dynamic magnetic field is applied to steer the FFP in the field of view. The movement of the FFP along a given trajectory constitutes a scan: the change of magnetization of the iron oxide particles due to the movement of the FFP induces a voltage in a set of receive coils which constitutes the measured data $f$. The relation between the concentration $u$ and the measured data $f$ is modeled as a linear mapping called the system function or system matrix $A$. In order to determine the system matrix/function both model and measurement-based approaches have been applied; cf. [48, 36] for instance. At present, the measurement-based approach is the most popularly used in real data scenarios: the system matrix is obtained by measuring the system responses to delta probes centered at a voxel. We consider MPI reconstruction in three dimesional space using real measurement data from the Open MPI dataset [39]. Hence, the reconstructed 3D image/volume $\tilde{u}$ lives in the space $X=\mathbb{R}^{n_{1}\times n_{2}\times n_{3}},$ where $n_{1}=n_{2}=n_{3}=19$ for the Open MPI data set. We denote the total number of voxels by $N=n_{1}n_{2}n_{3}=19^{3},$ and the number of measurements by $M$; here $M$ depends on the preprocessing applied to the MPI data (cf. Section 2.2). In particular, a scan in the Open MPI dataset contains the data coming from three different channels and the calibration data is Fourier transformed. This means that for each of the three channels the real and imaginary parts of the Fourier coefficients are considered and stacked, yielding a real-valued submatrix $A_{l}$, where $l=1,2,3$ is the channel index. Finally, the full matrix $A$ is formed by stacking the submatrices $A_{l}$ for each channel $l$ (for a detailed explanation we refer the interested reader to [31]). In order to avoid notation issues, we use the following conventions: we use the symbol $u\in\mathbb{R}^{N}$ to denote the (columnwise) vectorization of the spatial distribution $\tilde{u}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}.$ We then identify the system matrix with a matrix $\tilde{A}\in\mathbb{R}^{M\times N}$ such that $\tilde{A}\tilde{u}=Au$ for all volumes $\tilde{u}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$. We further identify the adjoint $A^{\ast}$ of $A$ with the matrix $\tilde{A}^{T}$. Using this notation, we have the relation
@@
where $\tilde{A}\in\mathbb{R}^{M\times N}$ denotes the system matrix, $u$ denotes the discrete nanoparticle concentration, $f$ the measured data, and the symbol $\eta$ denotes the noise on the measurements. The reconstruction task consists of determining $u$ in (1) given measurements $f$ (and calibration system matrix $A$). This is challenging due to noise and the ill-conditioning of $A$, and requires regularization.
2.2 The Preprocessing of the Data
We briefly discuss the preprocessing of the Open MPI data employed. It is inspired but not identical to the preprocessing proposed in [31]. For this reason, and since we discuss our scheme w.r.t. different levels of preprocessing in Section 3, we briefly explain the used method.
Preprocessing is applied to both calibration and phantom scan data. Since the calibration data are stored in the Fourier domain we Fourier transform the scan data (given as a time series). We further split the real and imaginary parts of the Fourier coefficients for both calibration and phantom data. Although each of the preprocessing steps may decrease the number of rows of $A$, we always use the same symbol $M$ to denote these numbers.
@@ 2.2 The Preprocessing of the Data
Due to analog filtering, the measurement data show a very high variance in the frequencies up to 75kHz. We theqrefore exclude frequencies lower than 80kHz. Further frequency selection avoiding frequencies with systematic strong noise or with low signal-to-noise ratio (SNR) has been considered in [31]. To reduce the number of variables at play and to avoid estimations of the SNR we consider no SNR-based thresholding in the preprocessing here.
@@
The noisy measured phantom data $f$ is of the form $f=f_{\mathrm{cl.}}+b_{f}+\eta_{f}$, with clean signal $f_{\mathrm{cl.}}$ containing the information on $u,$ the background signal $b_{f}$ produced by the drive field and perturbations in the measurement chain, and noise $\eta_{f}$ (assuming it is additive in nature). Accordingly, we decompose the system matrix as $A=A_{\mathrm{cl}}+b_{A}+\eta_{A},$ with clean signals $A_{\mathrm{cl}},$ background $b_{A}$ and noise $\eta_{A}.$ To exclude the background signals $b_{f},$ $b_{A}$, background measurements with no phantom in the scanner are performed after the scan and during the calibration procedure. To account for the background signal to be dynamic, it is repeated after every $19$th delta probe scan (after running through the $x$-axis) and a corresponding convex combination is subtracted. Each phantom and delta scan is performed $1000$ times and the average is taken as input for the reconstruction. By linearity, the averaging and subtraction is in fact carried out in the Fourier domain. We refer to [31] for more details.
@@
The low rank approximation is obtained by selecting $K\leq\min\{M,N\}$ and by performing a randomized Singular Value Decomposition (rSVD) [31, 26], yielding the triple of matrices $(\tilde{U}_{K},\tilde{\Sigma}_{K},\tilde{V}_{K})$ and considering for the reconstruction the transformed matrix $\tilde{U}_{K}^{\ast}A\in\mathbb{R}^{K\times N}$ and the transformed data $\tilde{f}=\tilde{U}_{K}^{\ast}f\in\mathbb{R}^{K}$. Because we aim at using the rSVD obtained we store $\tilde{V}_{K}$ and $\mathrm{diag}(\tilde{\Sigma}_{K})\in\mathbb{R}^{K}$ as well as $\tilde{U}_{K}^{\ast}A$ and $\tilde{U}_{K}^{\ast}f$.
2.3 ZeroShot-PnP: The Proposed Plug-and-Play Approach with a Zero-Shot Denoiser
We now describe the proposed ZeroShot-$\ell^{1}$-PnP approach and its variant, the ZeroShot-PnP approach. The regularized inversion of the ill-posed MPI reconstruction problem in Equation (1) can be formulated as a minimization problem of the form